OpenStudy (anonymous):

5x + 9 = 9(x + 1) - 4x I thought I had the answer, but realized what I had was wrong. Please help.

OpenStudy (anonymous):

how did you solve it?

OpenStudy (anonymous):

5x + 9 = 9x + 9 - 4x 5x + 9 = 5x + 9 - 9 = - 9 5x + 0 = 5x /5 /5 0 = x

OpenStudy (anonymous):

when you divide both sides by 5 you get x=x not 0=x

OpenStudy (anonymous):

O.K. where did I go wrong?

OpenStudy (anonymous):